A VC fund has a total committed capital of $100 million. The fund has a lifetime of 10 years, and a management fee of 2% per year on committed capital. We expect total exit proceeds to be $175 million, and the VC will earn a 20% carry, with a basis of all committed capital. (a) What is the total return on investments that the portfolio of target companies needs to earn, in order for the VC to begin earning carry? (b) What is the total return on investments that we expect the portfolio of target companies to yield? (c) How much in total fees do we expect that the VC will earn?
Answer #1

(a) Total Return on Investment= (175+100*20%*10)/(100)*100

=(375/100)*100

=375%

Annual Return on Investment will be=375/10 = 37.5%

(b) Return on Investment= (175)/(100+100*2%*10)*100

= (175/120)*100

= 1.4583*100

= 145.83%

Annual Return on Investment will be= 145.83/10 = 14.583%

(c) Total fees earned by VC= 100 (20%-2%) = 18 million